2) In an RLC circuit, ? = 10 Ω, ? = 0.5 ??, and ? = 6.0 ??, the current in the inductor is ?? = 12 ? sin(?? + 0.2). The circuit operates at half of the resonance frequency.

a) Find the current in the resistor, capacitor and for the entire circuit.

b) Find the voltage in the resistor, capacitor, and inductor.

c) Find the emf of the AC source.
Answer #1

R с L given RLC are in woon series of so current - is same across inductor, resistance and capacitor . v I,= Ic= IR IR = 12 sX = WL = 20 GL 21 (1.45x103) x (0.5X153) XL = 4.56 n Xc = I - = WC 27&c 18.29 27 (1AS XIV ) (6x156) Xc= 54.66 = IR XR VR - VoEmf= 14400+ (54.72 - 219.46) 2 Emf= 14400 + 27 145.85 | 12145.8 {mf = {mg = 112.89 Emd í 113 volt